A cat's displacement is 15 meters to the right in 7.0 seconds. If, at the start of the 7.0 seconds, the cat was moving at a velocity of 2.0 m/s left what was its final velocity?
Answer #1

u = 2 m/s
s = 15 m
t = 7 s

Using Kinematics

s = ut+0.5 at^{2}
15=2+7+0.5a * 72
a= 0.0408 m/s^{2}

v = u+at
u = 2 + 0.0408 * 7
u = 2.2856m/s